OpenStudy (anonymous):

Compare and contrast the following functions with repeat to symmetry, domain, and range In each case explain the reason for the differences. a) p(x) = x^pi

OpenStudy (anonymous):

b) P(x) = x^22/7 c) P(x) = x^25/8 d) P(x) = x^223/71

OpenStudy (anonymous):

@ganeshie8 @iambatman @greenglasses

OpenStudy (anonymous):

So, to find the symmetry, you set the equation to P(-x). if P(-x) = P(x) it has symmetry (around the y-axis)

OpenStudy (anonymous):

okay so i graphed them

OpenStudy (anonymous):

Alright, then you've already pretty much solved your question then.

OpenStudy (anonymous):

so thats for a)

OpenStudy (anonymous):

so thats for b

OpenStudy (anonymous):

so thats for c

OpenStudy (anonymous):

Basically, if the x is put to the power of an irrational numbers, it does not exist for negative numbers.

OpenStudy (anonymous):

and this is d

OpenStudy (anonymous):

so what are the symmetry, domain and range for each one?

OpenStudy (anonymous):

(when it's less than 1)

OpenStudy (anonymous):

So symmetry is whether it has symmetry around the x-axis or y-axis.

OpenStudy (anonymous):

You can see that of the four, only b does, right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

well and b

OpenStudy (anonymous):

Domain is the possible x-values. Two have xER (pretend that's that E-like symbol) but two only exist for values of x larger than 0.

OpenStudy (anonymous):

Range is the possible y-values. You can see that for all but one, the only ones that exist are 0 or larger.

OpenStudy (anonymous):

wait so now I'm confused. can you compare and contrast the domain and range and axis of symmetry for each one?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

You know that only one has symmetry, right?

OpenStudy (anonymous):

i think b has symmetry

OpenStudy (anonymous):

Yes. So that means that in terms of symmetry, there's not much to compare other than to say that only b has symmetry along the y-axis.

OpenStudy (anonymous):

Do you know how to dtermine domain and range?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Okay, look at graph 1. the x^pi one.

OpenStudy (anonymous):

What x-values have points on the graph?

OpenStudy (anonymous):

well to pos inf

OpenStudy (anonymous):

To positive infinity, from what value?

OpenStudy (anonymous):

domain: x>0

OpenStudy (anonymous):

Exactly. (including 0, but I assume you just used > because there's no equal to/greater than sign on the keyboard). So can you use that to figure out the domain for the other three graphs too?

OpenStudy (anonymous):

but it excludes 0

OpenStudy (anonymous):

No it doesn't. What's 0^(pi)?

OpenStudy (anonymous):

okay so can you check all the domains and ranges i write

OpenStudy (anonymous):

c) domain: all real numbers, range: all real numbers d) domain: all real numbers, range: all real numbers

OpenStudy (anonymous):

for a) domain: all real numbers except x>_0, range: all real numbers except y>_0 b) domain: all real numbers, range: all real numbers except y>_ 0

OpenStudy (anonymous):

for a) domain: x>_0, range: y>_0 b) domain: all real numbers, range: y>_ 0 c) domain: all real numbers, range: all real numbers d) domain: all real numbers, range: all real numbers

OpenStudy (anonymous):

a) d) and b) look good, but c)'s range is off.

OpenStudy (anonymous):

wait so can you retype everything cause I'm confused

OpenStudy (anonymous):

a) domain: x is greater than or equal to 0 range: y is greater than or equal to 0 b) domain: x is all real numbers range: y is greater than or equal to 0 c) domain: x is...? range: y is...? d) domain: x is all real numbers range: y is all real numbers

OpenStudy (anonymous):

c) is x^(25/8), right? Look at the graph.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so what's the domain and range?

OpenStudy (anonymous):

d: x>_0 and range: all real #

OpenStudy (anonymous):

So, negative y-values have points on the graph?

OpenStudy (anonymous):

so can you explain to me in each case(symmetry, domain, and range) what causes the differences specifically?

OpenStudy (anonymous):

i think my d and r are correct

OpenStudy (anonymous):

For c) the range is y is equal to or greater than 0.

OpenStudy (anonymous):

oh yea i understand

OpenStudy (anonymous):

so can you explain to me in each case(symmetry, domain, and range) what causes the differences specifically?

OpenStudy (anonymous):

Okay, let's put aside a) for now. Look at the other three. Why don't you put them in square root form?

OpenStudy (anonymous):

Sorry, by square root form I mean root form.

OpenStudy (anonymous):

I've got to go to sleep so I'll just tl;dr this alright? If you put c) in root form, you'll see than if x is negative, it'll be in an even root and to the power of an odd number, and the even root of a negative number to the power of an odd number is not real. however, if you put b) in root form, you'll see that it's both in an odd root and to the power of an even number, meaning that negative values of x are real. For d), same thing, just that since its greater than three, the equation resembles a cubic equation.